OpenStudy (anonymous):

A 30 kg occupant in a car has a chance of surviving a crash if the deceleration is not more than than 3 m/s. calculate the force on a 75 kg person decelerating at the same rate?

OpenStudy (anonymous):

\[f=ma\] \[30\times3 m/s^{2} \]\[f=90 n \] deceleration power and i stopped here i dont even know if it was right !

OpenStudy (anonymous):

Isn't it just asking for the force experienced by a 75kg person experiencing acceleration of 3m/s^2? \[\Large \begin{array}{l} F = ma\\ F = 75 \cdot 3 = {\rm{225N}} \end{array}\]

OpenStudy (anonymous):

well if it's like this then it's simple